OpenStudy (anonymous):

Huilan is 6 years older than Thomas. The sum of their ages is 92. What is Thomas's age?

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

but thanks for trying

OpenStudy (jadeishere):

okay so here is the equation n + (n + 6) = 92

OpenStudy (jacob902):

i dont want to give the answer alway heres and example Alvin is 7 years younger than Elga The sum of their ages is 87 What is Elga's age?

OpenStudy (jacob902):

If Alvin is 7 years younger than Elga, then Alvin + 7 = Elga. If the sum of their ages is 87, then Alvin + Elga = 87. Substitute Alvin + 7 in the previous equation to get Alvin + (Alvin + 7) = 87. Reduce that and get 2 Alvin = 80 which means that Alvin = 40. That means that Elga = 47.

OpenStudy (jadeishere):

Basically what Jacob said. :) If you try to use the equation n + (n + 6) = 92 n is Thomas' age what you have to do is this; 1) simplify the equation 2n + 6 = 92 2) isolate the variable 2n = 86 n = 86/2 then go form there >.<

OpenStudy (anonymous):

can u guys give each metals i can't give u both metals

OpenStudy (jacob902):

sure